Bob purchases 25,000 shares of common stock of Super Duper for $100,000. Exactly one year later when the stock is selling for $5 per share, Bob sells 5,000 shares for $25,000 and holds on to the remaining 20,000 shares. Which of the following statements is correct?

Group of answer choices

Bob has a realized capital gain of $4,000 and an unrealized capital gain of $1,000.

Bob has a realized capital gain of $5,000 and an unrealized capital gain of $20,000.

Bob has a $25,000 realized capital gain.

Bob has a $5,000 unrealized capital gain.
Answer #1

Realized capital gain =Number of shares*(Share one year later -Common Stock/Number of shares)
=5000*(5-100000/25000) =5000

Unrealized capital gain =Number of shares held on *(Share one year later -Common Stock/Number of shares)
=20000*(5-100000/25000) =20000

Option b is correct option